KOJO IS TWICE AS OLD AS ESI WHO IN TURN IS 5 YEAR OLDER THAN KWEKU. IF THEIR TOTAL AGE IS 47,HOW OLD IS EACH OF THEM


Answer by stanbon(75887)

KOJO IS TWICE AS OLD AS ESI WHO IN TURN IS 5 YEAR OLDER THAN KWEKU. IF THEIR TOTAL AGE IS 47,HOW OLD IS EACH OF THEM
-----------------
Let Kweku be "x" yrs. old.
Then ESI is "x+5" yrs old
And KOJO is 2(x+5) yrs old
------------------------
EQUATION:
x + x+5 + 2x+10 = 47
4x+15 = 47
4x = 32
x = 8 (Kweku's age)
x+5 = 13 (ESI's age)
2x+10 = 26 (Kojo's age)
==============================
Cheers,
Stan H.